# CMM-Math: A Chinese Multimodal Math Dataset To Evaluate and Enhance the Mathematics Reasoning of Large Multimodal Models

### v2

@@ Introduction
In this paper, we propose a math-specific LMM (Math-LMM) for multimodal math reasoning, as illustrated in Figure 4. To tackle the problems with multiple figures, we introduce a mixed instruction with the interleaved text and image inputs. Inspired by (Liu et al. 2023a), Math-LMM primarily comprises a Vision Encoder for encoding image information, an Adapter for modality alignment, and a Large Language Model (LLM) for mathematical reasoning. Moreover, to capture the abilities of math reasoning, we train our model using three stages: foundation pre-training aligns general visual information with LLMs, foundation fine-tuning learns general multimodal abilities, and mathematical fine-tuning uses instruction learning to learn multimodal mathematical reasoning.
@@
We translate the input sample as a mix instruction to train Math-LMM. The template instruction is “This is text. [IMAGE1]. This is text. … This is text. [IMAGE2]. [IMAGE3] … ”, where [IMAGE] is the image embedding after alignment. First, we input the images into the vision encoder (e.g., ViT (Dosovitskiy et al. 2021), DFN5B (Fang et al. 2023)) to obtain the image representation with a fixed dimension. The vision encoder is pre-trained on large-scale visual or multimodal datasets, which can embed the image effectively. Then, we obtain the image embedding [IMAGE] by aligning the image representation with LLM using an adapter. Common designs for the adapter include Cross-Attention (Gheini, Ren, and May 2021), Q-Former (Li et al. 2023), and MLP. We chose a simple but equally effective two-layer MLP with GELU functions for our Adapter. Finally, we input the mixed restructure with interleaved image-text inputs to LLMs (e.g., Qwen2-7B-Instruct (Bai et al. 2023a)).
@@
In the first foundational pre-training phase, we pre-train the adapter to align the image input with the LLM using large-scale general multimodal datasets with image descriptions. Particularly, we only update the parameters of the adapter and fix the parameters of LLM and encoder. Here, we select several datasets for adapter pre-training, including LLaVA-Pretrain and LLaVA-CC3M-Pretrain-595K (Liu et al. 2023a), and Mantis-Instruct (Jiang et al. 2024).
@@
In the foundational fine-tuning phase, we focus on learning task processing capability using instruction tuning. We train the parameters of both the Adapter and the LLM modules. We primarily use datasets involving general domain problems, such as ShareGPT-4o [^2] https://huggingface.co/datasets/OpenGVLab/ShareGPT-4o, MMDU (Liu et al. 2024b), llava-en-zh-300k (Liu et al. 2023a), and CogVLM-SFT-311K (Zhu et al. 2023; Liu et al. 2023a).
@@
In the third mathematical fine-tuning phase, we fine-tune the proposed Math-LMM for improved mathematical capabilities by adjusting both the Adapter and LLM modules. We train our model on mathematics-related datasets, including our proposed CMM-Math, GSM8K (Yu et al. 2023), competition_math (Hendrycks et al. 2021b), blossom-math-v4333 https://huggingface.co/datasets/Azure99/blossom-math-v4, Vietnamese-395k-meta-math-MetaMathQA-gg-translated444 https://huggingface.co/datasets/5CD-AI/Vietnamese-395k-meta-math-MetaMathQA-gg-translated, Vietnamese-meta-math-MetaMathQA-40K-gg-translated555 https://huggingface.co/datasets/5CD-AI/Vietnamese-meta-math-MetaMathQA-40K-gg-translated, and Vietnamese-microsoft-orca-math-word-problems-200k-gg-translated666 https://huggingface.co/datasets/5CD-AI/Vietnamese-microsoft-orca-math-word-problems-200k-gg-translated.
Experiments
Experimental Setups
@@
We evaluate a series of LMMs using CMM-MATH, including the current state-of-the-art open-source and closed-source models. Particularly, we select InternLM-XComposer2.5-VL (InternLM-VL) (Zhang et al. 2024a) and CogVLM2-llama3-Chinese-chat (CogVLM2) (Wang et al. 2023) for open-source models, and for closed-source models, we employ Qwen-VL (Bai et al. 2023a), Gemini (Team et al. 2024) and GPT-4o (OpenAI 2023).
Figure 5: The prompt for GPT-4o used for scoring.
@@
First, we evaluate the performance of typical LMMs on CMM-MATH. Then, we conduct experiments to verify the effectiveness of MATH-LMM over MATHVISTA (Lu et al. 2024) and MATH-V (Wang et al. 2024) datasets. In this paper, we employ the Accuracy and GPT-4o scores to measure the performance of our model and baselines. Particularly, we use accuracy for choice and yes-no problems, and use the GPT-4o score for fill-in-the-blank, and analysis problems. For the GPT-4o score, we use GPT-4o to calculate the scores of generated answers by giving the problems, solutions, answers, and generated responses. We designed a prompt, as shown in Figure 3, referring to (Zheng et al. 2023).
Table 4: Detail of Math-LMM parameters
Vision Encoder Adapter LLM Total
0.63B 0.02B 7.61B 8.26B
Models Size Overall Alg AnaG Ari CombG Comb Cnt Desc GrphT Log MetG SolG Stat TransG

### v3

@@ 4. Our Proposed Method
In this paper, we propose a math-specific LMM (Math-LMM) as a strong baseline for multimodal math reasoning, as illustrated in Figure 3. To tackle the problems with multiple figures, we introduce a mixed instruction with the interleaved text and image inputs. Inspired by LLaVa (Liu et al., 2024b), Math-LMM primarily comprises a Vision Encoder for encoding image information, an Adapter for modality alignment, and a Large Language Model (LLM) for mathematical reasoning. Moreover, to capture the abilities of math reasoning, we train our model using three stages: 1) the foundation pre-training stage aligns general visual information with LLMs via text-image pairs; 2) the foundation fine-tuning stage learns general multimodal abilities based on foundation instructions; and 3) the mathematical fine-tuning stage uses instruction learning based on math instructions to learn multimodal mathematical reasoning.
@@
We translate the input sample as a mix instruction to train Math-LMM. The template instruction is “This is text. [IMAGE1]. This is text. … This is text. [IMAGE2]. [IMAGE3] … ”, where [IMAGE] is the image embedding after alignment. First, we input the images into the vision encoder (e.g., ViT (Dosovitskiy et al., 2021), DFN-5B-H-14+ (Fang et al., 2023)) to obtain the image representation with a fixed dimension. The vision encoder is pre-trained on large-scale visual or multimodal datasets, which can embed the image effectively. Then, we obtain the image embedding [IMAGE] by aligning the image representation with LLM using an adapter. Common designs for the adapter include Cross-Attention (Gheini et al., 2021), Q-Former (Li et al., 2023), and MLP. We chose a simple but equally effective two-layer MLP with GELU functions for our Adapter. Finally, we input the mixed restructure with interleaved image-text inputs to LLMs (e.g., Qwen2-7B-Instruct (Bai et al., 2023a)).
@@
In the first foundational pre-training phase, we pre-train the adapter to align the image input with the LLM using large-scale general multimodal datasets with image descriptions. In particular, we only update the parameters of the adapter and fix the parameters of the LLM and encoder. Here, we select several datasets for adapter pre-training, including LLaVA-Pretrain and LLaVA-CC3M-Pretrain-595K (Liu et al., 2024b), and Mantis-Instruct (Jiang et al., 2024).
@@
In the foundational fine-tuning phase, we focus on learning task processing capability using instruction tuning. We train the parameters of both the Adapter and the LLM modules via large-scale foundation instructions. We primarily use datasets involving general domain problems, such as ShareGPT-4o [^2] https://huggingface.co/datasets/OpenGVLab/ShareGPT-4o, MMDU (Liu et al., 2024a), llava-en-zh-300k (Liu et al., 2024b), and CogVLM-SFT-311K (Zhu et al., 2023; Liu et al., 2024b).
@@
In the third mathematical fine-tuning phase, we fine-tune the proposed Math-LMM for improved mathematical capabilities by adjusting both the Adapter and LLM modules. We train our model on mathematics-related datasets, including our proposed CMM-Math, GSM8K (Cobbe et al., 2021), competition_math (Hendrycks et al., 2021b), blossom-math-v4333https://huggingface.co/datasets/Azure99/blossom-math-v4, Vietnamese-395k-meta-math-MetaMathQA-gg-translated , Vietnamese-meta-math-MetaMathQA-40K-gg-translated , and Vietnamese-microsoft-orca-math-word-problems-200k-gg-translated  [^4] https://huggingface.co/datasets/5CD-AI .
Figure 4. The prompt for GPT-4o used for scoring.
Table 4. Comparison of model performances in accuracy across different levels. The levels from 1 to 12 correspond to grades from primary to high school. The first and second highest accuracy of open-source LMMs are marked in red and blue.
Models Overall LV1 LV2 LV3 LV4 LV5 LV6 LV7 LV8 LV9 LV10 LV11 LV12
Open-source LMMs
CogVLM2 25.85 38.00 36.36 21.15 29.38 20.39 24.32 27.32 29.33 24.90 23.02 25.37 22.73
InternLM-VL 19.82 18.00 15.91 20.19 12.99 6.58 28.83 23.90 26.00 26.97 13.06 22.69 20.63
Qwen2-VL-Instruct 43.04 54.00 57.95 49.04 49.72 33.55 48.65 47.80 46.00 41.08 37.46 38.81 38.46
LLaVA-v1.5 18.08 40.00 27.27 27.88 21.47 17.76 25.23 15.12 18.67 18.67 15.81 12.84 9.09
LLaVA-v1.6-mistral 16.83 33.00 19.32 25.00 18.64 16.45 13.51 16.59 14.67 10.37 15.12 18.51 14.34
CogVLM2 (3-Shot) 31.21 37.00 34.09 40.38 29.94 28.95 35.14 34.63 32.67 28.22 28.52 32.84 25.52
InternLM-VL (3-Shot) 25.09 22.00 32.95 27.88 16.95 11.84 27.93 25.37 26.67 24.90 24.05 26.87 31.82
Qwen2-VL-Instruct (3-Shot) 46.29 38.00 62.50 53.85 55.93 54.61 60.36 47.80 46.00 44.81 40.89 36.42 43.01
LLaVA-v1.5 (3-Shot) 19.69 37.00 28.41 37.50 24.86 24.34 30.63 11.22 16.67 9.54 19.93 13.13 18.18
LLaVA-v1.6-mistral (3-Shot) 21.88 17.00 35.23 31.73 32.20 26.97 30.63 20.00 17.33 7.47 16.84 24.18 21.68
Math-LMM (Ours 7B) 32.10 32.00 44.32 37.50 35.59 32.24 35.14 32.20 28.67 32.37 30.58 31.94 26.22
Math-LMM (Ours 72B) 48.57 47.00 62.50 53.85 59.32 48.03 58.56 51.71 43.33 38.59 46.05 44.78 48.60
Closed-source LMMs
Qwen-VL-Max 49.91 70.00 62.50 56.73 54.80 46.71 67.57 46.83 43.33 41.08 45.70 48.96 46.85
Gemini 41.88 65.00 65.91 58.65 51.41 46.05 56.76 39.02 39.33 34.85 35.40 35.52 29.72
GPT-4o 29.02 46.00 43.18 38.46 33.90 32.89 32.43 27.32 24.67 28.63 19.59 27.76 23.78
Qwen-VL-Max (3-Shot) 64.91 60.00 76.14 72.12 69.49 57.24 71.17 64.39 58.67 52.70 66.67 68.06 67.83
Gemini (3-Shot) 41.65 69.00 64.77 52.88 41.24 46.05 55.86 38.54 41.33 39.83 33.33 33.73 34.97
GPT-4o (3-Shot) 65.98 80.00 84.09 75.00 81.92 69.74 72.07 68.78 61.33 56.85 52.23 66.57 59.44
Mean accuracy of LMMs 35.66 44.61 47.41 43.32 39.99 34.47 43.04 35.47 34.15 31.21 31.35 33.83 32.38
Table 5. Comparison of model performances in accuracy across various mathematical subjects. Alg: algebra, AnaG: analytic geometry, Ari: arithmetic, CombG: combinatorial geometry, Comb: combinatorics, Cnt: counting, Desc: deseriptive geometry, GrphT: graph theory, Log: logic, MetG: metric geometry, SolG: solid geometry, Stat: statistics, TransG: transformation geometry.
Models Overall Alg AnaG Ari CombG Comb Cnt Desc GrphT Log MetG SolG Stat TransG
Open-source LMMs
CogVLM2 25.85 21.94 25.22 24.76 28.85 24.68 26.54 29.79 44.44 19.05 33.79 23.36 45.45 25.58
InternLM-VL 19.82 21.94 26.38 20.68 21.15 12.99 16.67 19.15 0.00 9.52 20.48 12.15 27.27 13.95
Qwen2-VL-Instruct 43.04 36.45 42.03 50.65 36.54 40.26 45.68 27.66 44.44 39.68 43.34 37.85 45.45 37.21
LLaVA-v1.5 18.08 15.16 13.91 21.82 15.38 18.18 12.96 21.28 22.22 12.70 19.45 19.16 36.36 25.58
LLaVA-v1.6-mistral 16.83 17.10 16.52 22.31 7.69 11.69 10.49 8.51 0.00 19.05 13.99 15.42 27.27 16.28
CogVLM2 (3-Shot) 31.21 30.97 28.12 32.57 32.69 24.68 30.25 29.79 22.22 25.40 36.18 30.84 45.45 27.91
InternLM-VL (3-Shot) 25.09 29.03 30.14 24.76 15.38 27.27 25.93 12.77 0.00 26.98 24.57 17.76 9.09 25.58
Qwen2-VL-Instruct (3-Shot) 46.29 37.42 40.58 53.42 46.15 40.26 56.17 36.17 33.33 46.03 49.15 40.19 54.55 51.16
LLaVA-v1.5 (3-Shot) 19.69 19.03 14.20 28.66 3.85 23.38 21.60 14.89 11.11 15.87 14.68 16.36 45.45 2.33
LLaVA-v1.6-mistral (3-Shot) 21.88 22.26 17.97 32.25 5.77 20.78 22.22 4.26 11.11 28.57 15.70 13.55 27.27 16.28
Math-LMM (Ours 7B) 32.10 26.13 29.86 36.64 28.85 32.47 35.19 27.66 33.33 30.16 35.15 28.04 36.36 25.58
Math-LMM (Ours 72B) 48.57 48.06 44.35 60.59 32.69 50.65 53.70 17.02 22.22 57.14 44.03 35.98 63.64 27.91
Closed-source LMMs
Qwen-VL-Max 49.91 49.68 48.12 60.59 44.23 50.65 54.94 21.28 11.11 50.79 42.66 36.45 63.64 51.16
Gemini 41.88 36.77 37.97 55.37 26.92 37.66 48.15 31.91 33.33 26.98 37.54 32.71 54.55 25.58
GPT-4o 29.02 26.13 28.41 35.67 26.92 20.78 32.10 25.53 33.33 28.57 24.57 21.96 18.18 37.21
Qwen-VL-Max (3-Shot) 64.91 70.00 63.19 74.43 50.00 66.23 69.75 21.28 44.44 73.02 54.61 57.01 63.64 53.49
Gemini (3-Shot) 41.65 38.71 35.07 54.07 19.23 31.17 51.85 27.66 44.44 44.44 36.18 30.84 54.55 44.19
GPT-4o (3-Shot) 65.98 62.90 59.71 82.57 46.15 66.23 72.84 40.43 77.78 68.25 59.04 47.66 81.82 55.81
Mean accuracy of LMMs 35.66 33.87 33.43 42.88 27.14 33.33 38.17 23.17 27.16 34.57 33.62 28.74 44.44 31.27
Table 6. Comparison of model performances in GPT-4o score across different levels. The levels from 1 to 12 correspond to primary to high school grades. The maximum score is 10.
Models Overall LV1 LV2 LV3 LV4 LV5 LV6 LV7 LV8 LV9 LV10 LV11 LV12
Open-source LMMs
CogVLM2 2.82 2.96 3.66 3.52 2.99 2.29 2.90 2.46 3.05 2.37 2.43 2.45 2.43
InternLM-VL 4.48 4.05 5.54 5.27 4.90 4.20 5.12 4.00 4.03 2.60 4.13 4.41 4.35
Qwen2-VL-Instruct 5.26 4.23 5.54 6.07 5.40 5.18 6.05 5.51 5.53 4.21 4.99 4.94 4.64
LLaVA-v1.5 2.56 2.51 2.80 2.96 2.76 2.50 2.37 2.66 2.66 2.37 2.43 2.27 2.34
LLaVA-v1.6-mistral 2.81 2.55 2.88 2.94 2.57 2.43 2.42 2.91 3.36 3.85 3.04 2.68 2.59
CogVLM2 (3-Shot) 2.72 2.73 3.53 3.18 2.78 2.30 2.66 2.35 2.97 2.31 2.46 2.48 2.74
InternLM-VL (3-Shot) 4.35 3.95 5.46 5.06 4.62 3.77 4.78 3.96 4.05 3.08 4.11 4.31 4.36
Qwen2-VL-Instruct (3-Shot) 4.09 3.94 4.99 4.94 4.28 3.82 4.73 3.75 4.06 3.28 3.61 3.49 3.47
LLaVA-v1.5 (3-Shot) 3.34 2.75 2.75 3.10 2.70 3.14 2.49 3.77 4.89 5.79 3.75 2.88 3.42
LLaVA-v1.6-mistral (3-Shot) 3.78 3.27 3.28 3.50 3.30 3.86 2.62 3.78 5.67 6.23 4.31 3.05 3.73
Math-LMM (Ours 7B) 2.46 2.27 3.17 2.95 2.97 2.40 1.83 2.16 2.67 2.16 2.13 2.26 2.11
Math-LMM (Ours 72B) 4.04 3.61 4.66 4.52 3.71 3.97 4.49 3.81 3.84 3.64 4.00 3.93 3.86
Closed-source LMMs
Qwen-VL-Max 6.50 5.91 7.30 7.54 7.11 6.61 7.02 6.1 5.91 4.73 6.01 6.18 6.17
Gemini 6.02 6.30 6.85 6.83 6.46 5.75 6.40 5.67 6.46 5.01 5.42 4.96 5.62
GPT-4o 7.94 7.44 8.70 8.76 8.34 8.04 8.30 7.83 7.65 6.44 7.67 7.48 7.74
Qwen-VL-Max (3-Shot) 6.21 5.66 6.70 6.43 6.18 6.12 6.91 6.00 5.72 5.50 5.67 6.57 6.51
Gemini (3-Shot) 5.89 5.83 6.54 6.47 6.06 5.44 6.26 5.51 6.59 5.36 5.77 5.21 5.44
GPT-4o (3-Shot) 7.85 7.27 8.53 8.63 8.45 8.09 8.36 7.87 7.65 6.06 7.48 7.22 7.46
Mean accuracy of LMMs 4.62 4.29 5.16 5.15 4.75 4.44 4.76 4.45 4.82 4.17 4.41 4.26 4.39
Table 7. Comparison of model performances in GPT-4o score across various mathematical subjects.
Models Overall Alg AnaG Ari CombG Comb Cnt Desc GrphT Log MetG SolG Stat TransG
Open-source LMMs
CogVLM2 2.82 2.60 2.44 3.21 2.81 2.44 2.29 2.05 2.41 2.89 2.91 2.41 2.71 3.36
InternLM-VL 4.48 4.51 4.02 5.24 2.69 4.28 4.08 2.11 3.82 4.75 4.16 3.78 5.14 3.07
Qwen2-VL-Instruct 5.26 5.73 4.77 6.07 3.81 4.30 4.87 3.69 2.94 4.92 4.71 4.38 7.43 4.00
LLaVA-1.5 2.56 2.46 2.38 2.72 2.45 2.30 2.56 2.51 2.35 2.31 2.37 2.63 3.43 2.90
LLaVA-v1.6-mistral 2.81 2.47 3.35 2.68 2.73 2.60 3.09 3.57 3.71 2.52 2.93 2.61 3.86 3.74
CogVLM2 (3-Shot) 2.72 2.53 2.62 2.94 2.42 2.59 2.65 1.90 2.12 2.84 2.68 2.51 2.57 3.14
InternLM-VL (3-Shot) 4.35 4.58 4.14 5.08 2.91 4.25 3.73 2.06 3.06 4.44 3.91 3.65 4.71 2.93
Qwen2-VL-Instruct (3-Shot) 4.09 3.83 3.67 4.91 3.61 3.34 3.60 3.26 2.82 3.53 3.75 3.23 5.43 2.95
LLaVA-v1.5 (3-Shot) 3.34 2.91 3.92 2.74 5.36 2.76 3.23 5.27 3.82 2.95 4.04 3.86 3.43 4.90
LLaVA-v1.6-mistral (3-Shot) 3.78 3.27 4.07 2.94 5.90 3.08 3.51 6.24 4.82 3.27 5.05 5.02 2.43 4.83
Math-LMM (Ours 7B) 2.46 2.19 2.23 2.76 2.40 1.90 1.84 2.41 2.06 2.62 2.40 2.53 1.43 2.43
Math-LMM (Ours 72B) 4.04 4.25 3.90 4.52 2.80 3.59 3.56 3.26 2.65 4.44 3.74 3.46 7.00 3.24
Closed-source LMMs
Qwen-VL-Max 6.50 6.66 5.80 7.41 4.30 6.07 5.84 4.65 3.82 6.16 6.12 5.86 6.29 5.07
Gemini 6.02 5.65 5.29 6.72 5.72 5.15 5.32 5.17 5.47 5.30 6.06 5.71 6.00 5.05
GPT-4o 7.94 8.15 7.28 8.65 6.23 7.89 7.70 6.22 7.41 7.56 7.65 7.20 8.00 7.14
Qwen-VL-Max (3-Shot) 6.21 6.52 6.14 6.73 4.74 5.44 5.45 5.74 5.59 6.06 5.96 5.81 4.86 4.81
Gemini (3-Shot) 5.89 5.72 5.41 6.45 5.62 5.49 5.36 5.34 6.18 5.33 5.78 5.41 6.14 5.02
GPT-4o (3-Shot) 7.85 7.99 7.07 8.61 6.12 7.74 7.51 6.31 8.24 7.52 7.46 7.22 8.57 6.90
Mean accuracy of LMMs 4.62 4.56 4.36 5.02 4.03 4.18 4.23 3.99 4.07 4.41 4.54 4.29 4.97 4.19
Table 8. Accuracy scores on the testmini subset of MathVista. FQA: figure question answering, GPS: geometry problem solving, MWP: math word problem, TQA: textbook question answering, VQA: visual question answering. Mathematical reasoning types: ALG: algebraic reasoning, ARI: arithmetic reasoning, GEO: geometry reasoning, LOG: logical reasoning, NUM: numeric commonsense, SCI: scientific reasoning, STA: statistical reasoning.
Model Overall FQA GPS MWP TQA VQA ALG ARI GEO LOG NUM SCI STA
Open-source LMMs
IDEFICS-9B-Instruct 19.8 21.6 21.1 6.5 25.9 24.0 22.1 15.0 19.8 18.9 9.9 24.6 18.1
mPLUG-Owl-LLaMA-7B 22.2 22.7 23.6 10.2 27.2 27.9 23.6 19.2 23.9 13.5 12.7 26.3 21.4
miniGPT4-LLaMA-2-7B 23.1 18.6 26.0 13.4 30.4 30.2 28.1 21.0 24.7 16.2 16.7 25.4 17.9
LLaMA-Adapter-V2-7B 23.9 21.2 25.5 11.3 32.3 31.8 26.3 20.4 24.3 24.3 13.9 29.5 18.3
LLaVAR 25.2 21.9 25.0 16.7 34.8 30.7 24.2 22.1 23.0 13.5 15.3 42.6 21.9
InstructBLIP-Vicuna-7B 25.3 23.1 20.7 18.3 32.3 35.2 21.8 27.1 20.7 18.9 20.4 33.0 23.1
LLaVA-LLaMA-2-13B 26.1 26.8 29.3 16.1 32.3 26.3 27.3 20.1 28.8 24.3 18.3 37.3 25.1
Math-LMM (Ours 7B) 34.9 25.3 46.6 46.2 34.2 24.6 43.1 30.3 45.6 16.2 27.1 28.6 25.9
Math-LMM (Ours 72B) 36.3 36.4 33.7 46.2 36.1 29.1 29.9 34.6 33.5 24.3 28.5 41.8 40.2
Closed-source LMMs
Multimodal Bard 34.8 26.0 47.1 29.6 48.7 26.8 46.5 28.6 47.8 13.5 14.9 47.5 33.0
GPT-4V (Playground) 49.9 43.1 50.5 57.5 65.2 38.0 53.0 49.0 51.0 21.6 20.1 63.1 55.8
Human Performance
Human performance 60.3 59.7 48.4 73.0 63.2 55.9 50.9 59.2 51.4 40.7 53.8 64.9 63.9
5. Experiments
5.1. Experimental Setups
@@
We evaluate a series of LLMs on CMM-Math, including the current state-of-the-art open-source and closed-source models. Particularly, we select InternLM-XComposer2.5-VL (InternLM-VL) (Zhang et al., 2024a), Qwen2-VL-Instruct (Wang et al., 2024a) and CogVLM2-llama3-Chinese-chat (CogVLM2) (Wang et al., 2023) for open-source models, and for closed-source models, we employ Qwen-VL (Bai et al., 2023a), Gemini (Team et al., 2024) and GPT-4o (OpenAI, 2023).
@@
First, we evaluate the performance of typical LMMs on CMM-Math. Then, we conduct experiments to verify the effectiveness of MATH-LMM over MATHVISTA (Lu et al., 2024) and MATH-V (Wang et al., 2024b) datasets. In this paper, we employ the Accuracy and GPT-4o scores to measure the performance of our model and baselines. Particularly, we use accuracy for choice and yes-no problems, and use the GPT-4o score for fill-in-the-blank, and analysis problems. For the GPT-4o score, we use GPT-4o to calculate the scores of generated answers by giving the problems, solutions, answers, and generated responses. Specially, the prompt we designed is shown in Figure 3.
Implementation Details.
We employ a lightweight DFN5B-H-14+ (Fang et al., 2023) as the vision encoder. This model is trained on a large amount of high-quality paired image-text data using Contrastive Language Image Pre-training (CLIP) and demonstrates superior performance at a comparable scale of parameters. The image encoding module contains only 0.6 billion parameters, which enables our 7B model to perform inference deployment on a single 4090 GPU. In addition, we select an open-source, high-performance Qwen2 (Bai et al., 2023a) as our large language model module.
5.2. Evaluating SOTA LMMs on CMM-Math
We evaluate the performance of several SOTA LMMs over CMM-Math in terms of accuracy (Table 4 and Table 5) and GPT-4o score (Table 6 and Table 7). We also report the performance of various subjects and levels to analyze the LMMs’ abilities of multimodal mathematical reasoning. We give the analysis in two parts: attention to CMM-Math and attention to SOTA LMMs.
Attention to CMM-Math
We analyze CMM-Math from five perspectives: challenges of CMM-Math, comparisons across different subjects and levels, disparity between accuracy and GPT-4o score, and zero-shot versus few-shot scenarios.
Challenges of CMM-Math. From the last row of Table 4 and Table 5, we can observe that the average results across all models are only 35.66. In some subjects, the average results are even below 30, such as combinatorial, descriptive, solid geometry, and graph theory. Meanwhile, GPT-4o (3-Shot) achieves the best result with a score of 65.98 among all models. However, GPT-4o (3-Shot) still performs poorly on difficult levels, like Levels 9 and 10, and faces challenges in certain subjects, such as combinatorial, descriptive, and solid geometry. This indicates that achieving good results on CMM-Math is challenging for most current LMMs, and even the best LMMs struggle to perform exceptionally well on CMM-Math.
Comparisons Across Different Levels. Most LMMs perform well on primary school-level problems, but their performance declines at the high school level. As indicated in the last row of Table 4 and Table 6, the mean accuracy of LMMs shows an overall downward trend. Specifically, for instance, when employing three-shot prompting with GPT-4o, the model achieves an accuracy of 84.08 on level 2 but only 52.23 on level 10. This phenomenon can also be more intuitively observed in Figure 1, where for most models, the shaded area representing lower grade levels in the upper semicircle is significantly larger than the area representing higher grade levels in the lower semicircle.
Comparisons Across Different Subjects. Most LMMs excel in arithmetic and statistics but show limited proficiency in geometry. We can observe that the mean accuracy of all LMMs across all subjects is 35.66, with LMMs achieving an accuracy exceeding 40 in arithmetic and statistics subjects. In contrast, in geometry, including analytical, combinatorial, descriptive, metric, and transformation geometry, the accuracy of LMMs is below 35.66, with the worst performance at 23.17 for descriptive geometry.
Table 9. Comparison of model performances on MATH-V across various mathematical subjects. Alg: algebra, AnaG: analytic geometry, Ari: arithmetic, CombG: combinatorial geometry, Comb: combinatorics, Cnt: counting, DescG: descriptive geometry, GrphT: graph theory, Log: logic, Angle: metric geometry - angle, Area: metric geometry - area, Len: metric geometry-length, SolG: solid geometry, Stat: statistics, Topo: topology, TransG: transformation geometry.